OpenStudy (liddoislander):

Which of the following points is a solution of y < 3x - 1? A. (2, 0) B. (-2, 0) C. (0, 2) D. (-2, -2)

jimthompson5910 (jim_thompson5910):

I'll pick a point at random. I'll pick point D (-2,-2) That means x = -2 and y = -2 y < 3x - 1 -2 < 3(-2) - 1 ... plug in x = -2 and y = -2 -2 < -6 - 1 -2 < -7 The inequality -2 < -7 is false. It should be -2 > -7. So because -2 < -7 is false, this means y < 3x - 1 is false when (x,y) = (-2,-2) Therefore, (-2,-2) is NOT a solution to y < 3x - 1

jimthompson5910 (jim_thompson5910):

I'll let you check the others.

OpenStudy (liddoislander):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what makes you say C?